OpenStudy (anonymous):

30a. A company sells brass and steel machine parts. One shipment contains 3 brass and 10 steel parts and costs $48. A second shipment contains 7 brass and 4 steel parts and costs $54. Find the costs of each type of each type of machine parts. b. How much would a shipment containing 10 brass and 13 steel machine parts cost?

OpenStudy (anonymous):

x=brass y=steel 3x+10y=48 7x+4y=54 multiply first equation by -7 to get -21x-70y=-336 multiply second equation by 3 to get 21x+12y=162 combine -58y=-174 divide both sides by -58 y=3 put y in for first equation 3x+10(3)=48 3x+30=48 48-30=18 3x=18 divide both sides by 3 x=6

OpenStudy (anonymous):

is that a or b?

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

and since you only want the answer, $99 is b. merry christmas.

OpenStudy (anonymous):

i dont want just the answer i want to see how you solved it

OpenStudy (anonymous):

well you have the answer, try to work backwards! It'll get you the work you want, better yet, you'll learn :D

OpenStudy (anonymous):

can you just help me start it with the equation?

OpenStudy (anonymous):

@jollysailorbold ??

OpenStudy (anonymous):

since x=6 and y=3, 10x + 13y = 10(6) + 13(3) = 60 + 39 = 99 The shipment would cost $99.

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

im kind of confused on how that is all one equation

OpenStudy (anonymous):

nevermind i got it THANK YOU SO MUCH! YOU WERE SOO HELPFULL

OpenStudy (anonymous):

It's substitution. 10x + 13y you plug things in and get 10(6) + 13(3) and simplify to get 60 +39 and simplify further to get $99

OpenStudy (anonymous):

oh. haha okay :D pleasure!